OpenStudy (anonymous):

Jack is 4 years older than Jill. Seven years from now the sum of their ages will be 72. How old is jill now?

OpenStudy (anonymous):

76

OpenStudy (anonymous):

Divide 72 by 2. 72 / 2 = 36 Half of 4 is two, so add 2 to 36 for Jack's age seven years from now, and subtract 2 from 36 for Jill's age seven years from now. Jack will be 38 in seven years. Jill will be 34 in seven years. The question asks for Jill's age NOW. So subtract 7 from 34. 34 - 7 = 27. Jill is 27 now.

OpenStudy (anonymous):

@mayrabris2cute He said 76, not 26. But okay.

OpenStudy (anonymous):

Ok. Jill is 77. Solution: x+11+x+7= 72 2x+18=72 -18 2x=54 x= 23 years jill's age jack's age = 27

OpenStudy (anonymous):

Hmm, I'm trying to figure out how you got those numbers. How'd you get the x+11+x+7=72 ? I'm confused. Maybe I'm just tired. But I dunno how you did that. :V

OpenStudy (anonymous):

Thanks Eclipse. I totally understand now!

OpenStudy (anonymous):

Well. Idk how to explain it. I just got the equation! But @eclipse18 your prob right

OpenStudy (anonymous):

@mayrabris2cute Yeah I dunno how you got that, but thanks. x3 and @adair93 You're welcome, glad I could help!